OpenStudy (anonymous):

Find all solutions of the equation. 5 sin x - 8 sqrt2 = 3 sin x- 7 sqrt2

OpenStudy (anonymous):

A. x = 5 pi/4 + nπ or x =7 pi/4 + nπ B. x =pi/4 + nπ or x =3 pi/4 + nπ C. x = pi/4 + 2nπ or x = 3 pi/4 + 2nπ D. x = 5 pi/4 + 2nπ or x = 7 pi/4 + 2nπ

OpenStudy (anonymous):

@studystagram

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

get like terms together to simplify the equation

OpenStudy (anonymous):

oh hold on and let me see if I can work it out

OpenStudy (anonymous):

Ok i got x = pi/4 + 2nπ or x = 3 pi/4 + 2nπ all together is this correct?

OpenStudy (amistre64):

thats suits me just fine

OpenStudy (anonymous):

Thank you for the help @amistre64 :)

OpenStudy (studystagram):

x = pi/4 + 2npi or x = 3pi/4 + 2npi

OpenStudy (amistre64):

youre welcome :)

OpenStudy (anonymous):

and Thank you @studystagram